# Quantum automata and languages of finite index

### v2

@@ 6.1 The case of finite index matrix languages
Now we propose a first non trivial extension of the main results of decidability provided in [7]. Such an extension is based upon the machinery of Section 5, and some combinatorial tools developed by Choffrut, Varricchio and the second author in [10] for the study of rational relations. To this purpose, following [11], Ch. $1$, Sec. $1.1$, we first recall that a matrix context-free grammar (which we will henceforth simply call matrix grammar) is a tuple $G=\langle V,\Sigma,M,S\rangle$, where $V$, $\Sigma$, and $M$ are respectively the finite sets of nonterminals, terminals, and matrix (of rules), and $S\in V$ is the start symbol. Each matrix $m\in M$ is a finite sequence $m=(p_{1},\ldots,p_{s})$ where each $p_{i}$, $1\leq i\leq s$, is a context-free production from $V$ to $(V\cup\Sigma)^{*}$. We denote by $P_{M}$ the set of all the context-free productions that appear in the matrices of $M$. For an arbitrary $m\in M$, we define the $1$-step derivation in $G$, $x\mathrel{\mathop{\Rightarrow}_{m}}y,$ with $x,y\in(V\cup\Sigma)^{*}$ if
@@
where $\mathrel{\mathop{\Rightarrow}_{p_{i}}}$ is the standard context-free derivation relation. This is equivalent to say that $x=x_{0},x_{s}=y$, and, for every $i=1,\ldots,s,$
@@
where $p_{i}=(X_{i}\rightarrow\alpha_{i})$, and for some $x^{\prime}_{i-1},x^{\prime\prime}_{i-1}\in(V\cup\Sigma)^{*}$. In other words, a $1$-step derivation in the matrix grammar $G$ corresponds to a $s$-step derivation in the context-free grammar $\langle V,\Sigma,P_{M},S\rangle$, i.e., using, one by one, all the context-free productions of $m$. Let $M^{*}$ be the set of the finite sequences of matrices and $\alpha\in M^{*}$. If $\alpha=\varepsilon$, then $x\Rightarrow_{\alpha}x,x\in V^{*}$; if one has
@@
The following definition is instrumental.
@@
  • $(X_{1}\rightarrow u_{1}Y_{1}v_{1},\ldots,X_{k}\rightarrow u_{k}Y_{k}v_{k})$, for some $X_{i},Y_{i}\in V_{i},\ u_{i},\ v_{i}\in\Sigma^{*}$, with $1\leq i\leq k$.
  • For every $X_{i}\in V_{i}$ and $1\leq i\leq k$, $(X_{1}\rightarrow\varepsilon,\ldots,X_{k}\rightarrow\varepsilon)$.
@@
Let $L=\{a^{n}b^{n}c^{n}:n\in\mathbb{N}\}$ be the language over the alphabet $\Sigma=\{a,b,c\}$. Let $G$ be the restricted matrix grammar with the matrices of rules $m_{1}=(S\rightarrow ABC)$, $m_{2}=(A\rightarrow aA,B\rightarrow bB,C\rightarrow cC),$ and $m_{3}=(A\rightarrow~{}\varepsilon,B\rightarrow~{}\varepsilon,C\rightarrow\varepsilon).$ Note that the other components of $G$ can be deduced from these matrices. One easily checks that $L=L(G)$. Indeed, each word $a^{n}b^{n}c^{n},n\geq 0,$ can be generated by using first the matrix $m_{1}$, then the matrix $m_{2}$ $n-1$ times and finally the matrix $m_{3}$. Hence $L\subseteq L(G)$. The reverse inclusion is proved similarly.
@@
Let $L$ be a language generated by a restricted grammar and let $Q$ be a quantum automaton. Then one has:
@@
(i) Set ${\cal X}=(V^{\prime})^{k}$ and ${\cal Y}=\{X_{1}\cdots X_{k}\in{\cal X}:\exists\ m\in M:m=(S\rightarrow X_{1}\cdots X_{k})\}$. It is easily checked that, given $w\in\Sigma^{*}$, $w\in L$ if and only if there exists some $\omega=(u_{1},v_{1},u_{2},v_{2},\ldots,u_{k},v_{k})\in(\Sigma^{*})^{2k},$ such that
@@
Let $s=(q_{1},\ldots,q_{n})$ be a repetition-free sequence of elements of $\cal X$, ie, for every positive integers $i,j$, if $i\neq j$, then $q_{i}\neq q_{j}$. Clearly, $n$ is less than or equal to the number of states of $\cal X$ and therefore the set $\cal S$ of all repetition-free sequences is finite. Further, for every $p,q\in\cal X$, let ${\cal S}_{pq}$ be the set
@@
Now, for every $p,q\in\cal X$, define the subset $E_{pq}$ of $(\Sigma^{*})^{2k}$ as:
@@
and the subset $C_{q}$ of $(\Sigma^{*})^{2k}$ as:
$C_{q}=\{\;(u_{1},v_{1},\ldots,u_{k},v_{k})\;:\;\exists\ \alpha\in M^{*}:\ X_{1}\cdots X_{k}\Rightarrow_{\alpha}u_{1}Y_{1}v_{1}\cdots u_{k}Y_{k}v_{k}\;\}.$
One easily checks that $C_{q}$ is a submonoid of the monoid $(\Sigma^{*})^{2k}$. For every $s=(q_{1},\ldots,q_{n})\in\cal S$, define the (possibly empty) subset ${\cal L}_{s}$ of $(\Sigma^{*})^{2k}$ as:
${\cal L}_{s}=C_{q_{1}}E_{q_{1}q_{2}}C_{q_{2}}\cdots C_{q_{n-1}}E_{q_{n-1}q_{n}}C_{q_{n}}$ (28)
By using the very same argument of [10], Theorem 2, one can prove that
${\cal L}_{pq}=\bigcup_{s\in{\cal S}_{pq}}\ {\cal L}_{s}$ (29)
Let now consider the morphism
@@
$\widehat{\varphi}(\omega)=\varphi(u_{1})\ \oplus\ \varphi(v_{1})\ \oplus\ \cdots\ \oplus\ \varphi(u_{k})\ \oplus\ \varphi(v_{k}),$
of the $2k$-fold product $O_{2nk}$ of $O_{n}.$
Now we prove that ${\mathbf{C}l}({\varphi}(L))$ is semialgebraic. By Proposition 3, it is enough to show that ${\mathbf{C}l}(\widehat{\varphi}(\cal L))$ is so. To this purpose, with an arbitrary nonterminal $q\in{\cal X}$ of $G$, associate the subset $M_{q}$ of $O_{2nk}$ defined as
$\{\varphi(u_{1})\oplus\varphi(v_{1})^{\scriptstyle T}\oplus\cdots\oplus\varphi(u_{k})\ \oplus\ \varphi(v_{k})^{\scriptstyle T}\mid(u_{1},v_{1},\ldots,u_{k},v_{k})\in C_{q}\}$ (30)
By using the very same argument of Lemma 5, one proves that $M_{q}$ is a monoid and its closure ${\mathbf{C}l}(M_{q})$ is an algebraic group. From this, by Proposition 1, it follows that ${\mathbf{C}l}(\widehat{\varphi}({C}_{q}))$ is algebraic as well. This implies, via Corollary 1 and Proposition 3, that every set ${\mathbf{C}l}(\widehat{\varphi}({\cal L}_{s}))$ of (28) is semialgebraic. Then the semialgebraicity of ${\mathbf{C}l}(\widehat{\varphi}(\cal L))$ follows, by applying Proposition 1, from the last condition and (27), (29).
(ii) Assume that the quantum automaton ${\cal Q}$ is rational. First we prove that ${\mathbf{C}l}(M_{q}))$ is effective algebraic. One proceeds as in the proof of Lemma 6. Indeed, since $L$ is generated by a restricted matrix grammar, then $M_{q}$ is a regular submonoid of $O_{2nk}$. Precisely, it is recognized by the finite $O_{2nk}$–automaton whose states are the elements of $\cal X$, the initial and final states coincide with $q$, and the transitions are of the form
@@
Hence $M_{q}$ is regular and the group $\langle M_{q}\rangle$ has an effective finite generating set. By applying to $\langle M_{q}\rangle$ the algorithm of Derksen et al. [12], one gets an effective presentation of ${\mathbf{C}l}(M_{q})$. From this, by Proposition 1, it follows that ${\mathbf{C}l}(\widehat{\varphi}({C}_{q}))$ is effective algebraic as well. Finally, since all the steps of the proof (i) are effective, the fact that ${\mathbf{C}l}(\widehat{\varphi}({C}_{q}))$ is effective algebraic implies that ${\mathbf{C}l}({\varphi}(L))$ is so. The claim follows by applying Proposition 7. ∎∎
We now discuss some straightforward corollaries of Proposition 10. First, remark that, by applying the proposition to the matrix grammars of index $k=1$, we obtains [7], Prop. 22 (cf Corollary 2). Moreover, the language of Example 2 is an instance of the following more general result. To this purpose, we recall that a bounded semilinear language is a finite union of linear languages which are languages of the form
$L=\{u_{1}^{n_{1}}\cdots u_{k}^{n_{k}}\mid(n_{1},\ldots,n_{k})\in{R}\}$ (31)
@@
$R=\{{\bf v}_{0}+\lambda_{1}{\bf v}_{1}+\cdots+\lambda_{\ell}{\bf v}_{\ell}\mid\lambda_{1},\ldots,\lambda_{\ell}\in\mathbb{N}\}$
@@
It is enough to prove the claim for an arbitrary language $L$ of the form (31). It is checked that $L$ is generated by the restricted grammar $G=\langle V,\Sigma,M,S\rangle$, where $V=\{S,X_{1},\ldots,X_{k}\}$ and $M$ is given by the matrix rules
  • $(S\rightarrow X_{1}\cdots X_{k})$.
  • $(S\rightarrow u_{1}^{{\bf v}_{01}}X_{1}\cdots u_{k}^{{\bf v}_{0k}}X_{k})$, with ${{\bf v}_{0}=({\bf v}_{01}},\ldots,{\bf v}_{0k})$.
@@
Example 2 (continued). Let ${\cal Q}=(s,\varphi,P,\lambda)$ be a finite quantum automaton where $\varphi:\Sigma^{*}\longrightarrow O_{n}$ is the morphism (4) associated with ${\cal Q}$. Recall that $\mathbf{C}l{({\cal A}))}$ and $\overline{{\cal A}}$ denote, respectively, the Euclidean and the Zariski closure of a set $\cal A$. The set ${\mathbf{C}l}(\varphi(L))$ is then equal to
@@
We finally observe that Proposition 10 is a non trivial extension of Corollary 2 and Corollary 4. Indeed the following example provides languages, neither bounded nor context-free, defined by restricted grammars.
@@ 6.2 The case of monoidal languages
We start with the following definition that is instrumental for this section.
@@
of $k$ families of linear grammars where the productions of $G$ satisfy the following conditions:
  • if $\ (X\rightarrow u)\in P,\,\ u\in\Sigma^{*}$ (terminal production)  then $u=\varepsilon$
[…]
@@
Let $(L,Q)$ be a pair where $Q$ is a rational quantum automaton and $L$ is a language generated by a monoidal grammar $G={\cal G}_{1}\circ{\cal G}_{2}\circ\cdots\circ{\cal G}_{k}$. If the Zariski closures of the monoids of cycles of the grammars of ${\cal G}_{k}$ are irreducible, then ${\mathbf{C}l}(\varphi(L))$ is effective semialgebraic.
We will do the proof of Proposition 11 for $k\leq 2$, i.e., for a monoidal grammar $G={\cal G}_{1}\circ{\cal G}_{2}$ of index $2$, since, by (17), the general case is treated with the same argument. The proof is technically structured in the following lemmata.
Now we find useful to recall a notation introduced in Section 2.4.
@@
Let $G={\cal G}_{1}$ be a monoidal grammar of index $1$ and let $L=L(G)$. If the Zariski closure of all the monoids of cycles of $G$ are irreducible, effective algebraic sets, then there exist a set $H\in({\cal H})$ and a regular map $\Psi$ such that
${\mathbf{C}l}(\varphi(L))=\Psi(H).$
In particular ${\mathbf{C}l}(\varphi(L))$ is effective semialgebraic and contains $I$.
@@
Let $G=\langle V,\Sigma,P,S\rangle$ be a monoidal grammar of index $1$. We rerun the proof of Proposition 9 for $G$. We proceed thus on $\mathop{\mathrm{Card}}\nolimits(V)$. If $\mathop{\mathrm{Card}}\nolimits(V)=1,$ i.e., $V=\{S\}$, then $G$ has a sole monoid of cycles $M_{S}$. Taking into account (26) and the form of the terminal productions of $G$, one has ${\mathbf{C}l}(\varphi(L_{S}))=\Psi({\mathbf{C}l}(M_{S}))$, where $\Psi(X\oplus Y):=XY^{\scriptstyle T}.$ From this and the fact that, by hypothesis, ${\mathbf{C}l}(M_{S})$ is irreducible, it follows the claim in this case. The basis of induction is proved.
Let us prove the inductive step. By Proposition 8, and taking into account the form of the productions of $G$, one has
$L_{S}=\bigcup_{i=1}^{m}\,C_{S}\diamond L_{A_{i}},$ (34)
where, for every $i=1,\ldots,m$, $A_{i}\in V^{\prime}=V\setminus\{S\}$. Relation (34) now implies
${\mathbf{C}l}(\varphi(L_{S}))=\bigcup_{i=1}^{m}\,{\mathbf{C}l}(M_{S}\diamond\varphi(L_{A_{i}})).$
By the induction hypothesis, for every $i=1,\ldots,m,$ ${\mathbf{C}l}(\varphi(L_{A_{i}}))=\Psi^{\prime}(H_{i})$, where $H_{i}\in({\cal H})$, $\Psi^{\prime}$ is a regular map (not depending on the variable $A_{i}$) and ${\mathbf{C}l}(\varphi(L_{A_{i}}))$ contains $I$. From this, by a straightforward computation, one gets
${\mathbf{C}l}(\varphi(L_{S}))=\Psi^{\prime\prime}(\,\bigcup_{i=1}^{m}\,{\mathbf{C}l}(M_{S})\oplus H_{i}\,),$ (35)
where
$\Psi^{\prime\prime}(X\oplus Y\oplus Z):=X\Psi^{\prime}(Z)Y^{\scriptstyle T}.$
Since ${\mathbf{C}l}(M_{S})$ is effective algebraic and $H_{i}\in({\cal H}),1\leq i\leq m$, then, by Proposition 5, the set $\bigcup_{i=1}^{m}\,{\mathbf{C}l}(M_{S})\oplus H_{i}\in({\cal H})$ as well. Finally the claim comes from (35) taking into account that the map $\Psi^{\prime\prime}$ is regular. ∎∎
Remark 3
Note that the map $\Psi$ of the statement of Lemma 8 does not depend upon the variable used as axiom of $G$.
Let ${\cal G}_{1}=\langle V_{1},\Sigma_{1},P_{1},S\rangle$ be the first grammar of the composition $G={\cal G}_{1}\circ{\cal G}_{2}$. Let $\Sigma_{1}^{*}\times\Sigma_{1}^{*}$ be the multiplicative monoid of binary relations on $\Sigma_{1}^{*}$ and let $\Pi(O_{n}\oplus O_{n})$ be the multiplicative monoid of subsets of $O_{n}\oplus O_{n}$. Define now the morphism between the two monoids
@@
For every $S\in V_{1}$, there exists an effective computable regular subset $\cal R$ of $\ \Sigma_{1}^{*}\times\Sigma_{1}^{*}$ such that $M_{S}=\zeta({\cal R})$.
@@
Finally observe that, by (38), $(\alpha,\beta)\in C_{S}$ so that $(\alpha,\beta^{\sim})\in\widehat{C}_{S}$. By (40) the latter implies $\varphi(u)\oplus\varphi(v)^{\scriptstyle T}\in\zeta(\widehat{C}_{S})$. Hence $M_{S}\subseteq\zeta(\widehat{C}_{S})$. The other inclusion is checked similarly.
Since ${\cal G}_{1}$ is a linear grammar, by a well-known construction (cf [4]), one has that $\widehat{C}_{S}$ is an effective computable rational binary relation on $\Sigma_{1}^{*}$. The lemma is proved. ∎∎
@@
For every $A\in V_{1}$, ${\mathbf{C}l}(M_{A})\in({\cal H}).$ In particular, ${\mathbf{C}l}(M_{A})$ is effective algebraic.

### v3

@@ 6.1 The case of finite index matrix languages
Now we propose a first non trivial extension of the main results of decidability provided in [10]. Such an extension is based upon the machinery of Section 5, and some combinatorial tools developed by Choffrut, Varricchio and the second author in [13] for the study of the counting functions of rational relations. To this purpose, following [14], Ch. $1$, Sec. $1.1$, we first recall that a matrix context-free grammar (which we will henceforth simply call matrix grammar) is a tuple $G=\langle V,\Sigma,M,S\rangle$, where $V$, $\Sigma$, and $M$ are respectively the finite sets of nonterminals, terminals, and matrices (of rules), and $S\in V$ is the start symbol. Each matrix $m\in M$ is a finite sequence $m=(p_{1},\ldots,p_{s})$ where each $p_{i}$, $1\leq i\leq s$, is a context-free production from $V$ to $(V\cup\Sigma)^{*}$. We denote by $P_{M}$ the set of all the context-free productions that appear in the matrices of $M$. For an arbitrary $m\in M$, we define the $1$-step derivation in $G$, $x\mathrel{\mathop{\Rightarrow}_{m}}y,$ with $x,y\in(V\cup\Sigma)^{*}$ if
@@
where $\mathrel{\mathop{\Rightarrow}_{p_{i}}}$ denotes the standard context-free derivation relation, defined by the rule $p_{i}$. This is equivalent to say that $x=x_{0},x_{s}=y$, and, for every $i=1,\ldots,s,$
@@
where $p_{i}=(X_{i}\rightarrow\alpha_{i})$, and for some $x^{\prime}_{i-1},x^{\prime\prime}_{i-1}\in(V\cup\Sigma)^{*}$. In other words, a $1$-step derivation in the matrix grammar $G$ corresponds to a $s$-step derivation in the context-free grammar $\langle V,\Sigma,P_{M},S\rangle$, i.e., using, one by one, all the context-free productions of $m$ (w.r.t. the order the $p_{i}$’s appear in $m$). Let $M^{*}$ be the set of the finite sequences of matrices and $\alpha\in M^{*}$. If $\alpha=\varepsilon$, then $x\Rightarrow_{\alpha}x,x\in V^{*}$; if one has
@@
The following definition is instrumental. It corresponds to a specific case of the simple matrix grammars introduced by Ibarra in [24] (see also [27]).
@@
  • $(X_{1}\rightarrow u_{1}Y_{1}v_{1},\ldots,X_{k}\rightarrow u_{k}Y_{k}v_{k})$, for some $X_{i},Y_{i}\in V_{i},\ u_{i},\ v_{i}\in\Sigma^{*}$, with $1\leq i\leq k$,
  • $(X_{1}\rightarrow\varepsilon,\ldots,X_{k}\rightarrow\varepsilon)$, with $X_{i}\in V_{i},$ $1\leq i\leq k$,
@@
Such grammars are very akin to simple matrix linear grammars (cf [24], pag. 363, [14], Ch. $1$, Sec. $5$).
@@
Let $L=\{a^{n}b^{n}c^{n}:n\in\mathbb{N}\}$ be the language over the alphabet $\Sigma=\{a,b,c\}$. Let $G$ be the restricted matrix grammar with the matrices of rules $m_{1}=(S\rightarrow ABC)$, $m_{2}=(A\rightarrow aA,B\rightarrow bB,C\rightarrow cC),$ and $m_{3}=(A\rightarrow~{}\varepsilon,B\rightarrow~{}\varepsilon,C\rightarrow\varepsilon).$ Note that the other components of $G$ can be deduced from these matrices. One easily checks that $L=L(G)$. Indeed, each word $a^{n}b^{n}c^{n},n\geq 0,$ can be generated by using first the matrix $m_{1}$, then the matrix $m_{2}$ $n$ times and finally the matrix $m_{3}$. Hence $L\subseteq L(G)$. The reverse inclusion is proved similarly.
@@
Let $L$ be a language generated by a restricted grammar and let $\cal Q$ be a quantum automaton. Then one has:
@@
Let $L$ be a language generated by a grammar as in Definition 9. Let us prove (i). Set ${\cal X}=(V^{\prime})^{k}$, and consider the sets
${\cal Y}=\{X_{1}\cdots X_{k}\in{\cal X}:\exists\ m\in M:m=(S\rightarrow X_{1}\cdots X_{k})\},$
and
${\cal Z}=\{X_{1}\cdots X_{k}\in{\cal X}:\exists\ m\in M:m=(X_{1}\rightarrow\varepsilon,\ldots,X_{k}\rightarrow\varepsilon)\}.$
Denote by $(\Sigma^{*})^{2k}$ the multiplicative monoid of $2k$-fold relations on $\Sigma^{*}$. It is easily checked that, given $w\in\Sigma^{*}$, $w\in L$ if and only if there exists some $\omega=(u_{1},v_{1},u_{2},v_{2},\ldots,u_{k},v_{k})\in(\Sigma^{*})^{2k},$ such that
@@
Let $s=(q_{1},\ldots,q_{n})$ be a repetition-free sequence of elements of $\cal X$, i.e., for every positive integers $i,j$, if $i\neq j$, then $q_{i}\neq q_{j}$. Clearly, $n$ is less than or equal to the cardinal number of $\cal X$. Denote by $\cal S$ the set of all such repetition-free sequences and observe that this set is finite. Further, for every $p,q\in\cal X$, let ${\cal S}_{pq}$ be the set
@@
Now, for every $p=X_{1}\cdots X_{k},q=Y_{1}\cdots Y_{k}$, define the subset $E_{pq}$ of $(\Sigma^{*})^{2k}$ as:
@@
and the subset $C_{p}$ of $(\Sigma^{*})^{2k}$ as:
$C_{p}=\{\;(u_{1},v_{1},\ldots,u_{k},v_{k})\;:\;\exists\ \alpha\in M^{*}:\ X_{1}\cdots X_{k}\Rightarrow_{\alpha}u_{1}X_{1}v_{1}\cdots u_{k}X_{k}v_{k}\;\}.$
For every $s=(q_{1},\ldots,q_{n})\in\cal S$, define the (possibly empty) subset ${\cal L}_{s}$ of $(\Sigma^{*})^{2k}$ as:
${\cal L}_{s}=C_{q_{1}}\diamond E_{q_{1}q_{2}}\diamond C_{q_{2}}\diamond\cdots\diamond C_{q_{n-1}}\diamond E_{q_{n-1}q_{n}}\diamond C_{q_{n}},$ (27)
where ()($\diamond$)( ⋄ ) corresponds to the extension of the operation (15) to $(\Sigma^{*})^{2k}$, that is, for subsets ${\cal M},{\cal M}^{\prime}$ of $(\Sigma^{*})^{2k}$, the tuple $(w_{1},\ldots,w_{2k})\in{\cal M}\diamond{\cal M}^{\prime}$ if and only if there exist
$(u_{1},v_{1},\ldots,u_{k},v_{k})\in{\cal M}\;\;\mbox{ and }\;\;(u^{\prime}_{1},v^{\prime}_{1},\ldots,u^{\prime}_{k},v^{\prime}_{k})\in{\cal M}^{\prime}$
such that, for every $i=1,\ldots,k,$
$w_{2i-1}\ =\ u_{i}u^{\prime}_{i}\;\;\;\;\mbox{ and }\;\;\;\;w_{2i}=v^{\prime}_{i}v_{i}.$
By using an argument similar to that of [13], Theorem 2, one checks that
${\cal L}_{pq}=\bigcup_{s\in{\cal S}_{pq}}\ {\cal L}_{s}.$ (28)
Let now consider the monoid morphism
@@
$\widehat{\varphi}(\omega)=\varphi(u_{1})\ \oplus\ \varphi(v_{1})\ \oplus\ \cdots\ \oplus\ \varphi(u_{k})\ \oplus\ \varphi(v_{k}).$
Since, by Theorem 1, one has
${\mathbf{C}l}({\varphi}(L))\ =\ \{M_{1}\cdots M_{2k}\ :M_{1}\oplus\cdots\oplus M_{2k}\in{\mathbf{C}l}(\widehat{\varphi}(\cal L))\},$
to show that ${\mathbf{C}l}({\varphi}(L))$ is semialgebraic, by Proposition 4, it is enough to show that ${\mathbf{C}l}(\widehat{\varphi}(\cal L))$ is so. To this purpose, with an arbitrary $q\in{\cal X}$, associate the subset $M_{q}$ of $O_{2nk}$ defined as
$\{\varphi(u_{1})\oplus\varphi(v_{1})^{\scriptstyle T}\oplus\cdots\oplus\varphi(u_{k})\ \oplus\ \varphi(v_{k})^{\scriptstyle T}\mid(u_{1},v_{1},\ldots,u_{k},v_{k})\in C_{q}\}.$ (29)
By proceeding as in Lemma 3, one proves that $M_{q}$ is a monoid and its closure ${\mathbf{C}l}(M_{q})$ is an algebraic group. From this, by Proposition 1, it follows that ${\mathbf{C}l}(\widehat{\varphi}({C}_{q}))$ is an algebraic set. This implies, via Theorem 1 and Proposition 2, that every set ${\mathbf{C}l}(\widehat{\varphi}({\cal L}_{s}))$ of (27) is semialgebraic. Then the semialgebraicity of ${\mathbf{C}l}(\widehat{\varphi}(\cal L))$ follows, by applying Proposition 1, from the last condition and (26), (28).
(ii) Assume that the quantum automaton ${\cal Q}$ is rational. First we prove that ${\mathbf{C}l}(M_{q})$ is effective algebraic. One proceeds as in the proof of Lemma 4. Indeed, since $L$ is generated by a restricted matrix grammar, then $M_{q}$ is a regular submonoid of $O_{2nk}$. Precisely, it is recognized by the finite $O_{2nk}$–automaton whose states are the elements of $\cal X$, the initial and final states coincide with $q$, and the transitions are of the form
@@
Hence $M_{q}$ is a regular monoid and the group $\langle M_{q}\rangle$ has an effective finite generating set. By applying to $\langle M_{q}\rangle$ the algorithm of Derksen et al. one gets an effective presentation of ${\mathbf{C}l}(M_{q})$. From this, by Proposition 1, it follows that ${\mathbf{C}l}(\widehat{\varphi}({C}_{q}))$ is effective algebraic as well. Finally, since all the steps of the proof (i) are effective, the fact that ${\mathbf{C}l}(\widehat{\varphi}({C}_{q}))$ is effective algebraic implies that ${\mathbf{C}l}({\varphi}(L))$ is effective semialgebraic as well. The claim follows by applying Proposition 7. ∎∎
We now discuss some straightforward corollaries of Proposition 10. First, remark that, by applying the proposition to the matrix grammars of index $k=1$, we recover [10], Prop. 22 (cf Corollary 2). Moreover, note that the language of Example 2 is an instance of the following more general result. To this purpose, we recall that a bounded semilinear language is a finite union of linear languages which are languages of the form
$L=\{u_{1}^{n_{1}}\cdots u_{k}^{n_{k}}\mid(n_{1},\ldots,n_{k})\in{R}\},$ (30)
@@
$R=\{{\bf v}_{0}+\lambda_{1}{\bf v}_{1}+\cdots+\lambda_{\ell}{\bf v}_{\ell}\mid\lambda_{1},\ldots,\lambda_{\ell}\in\mathbb{N}\}.$ (31)
@@
It is enough to prove the claim for an arbitrary language $L$ of the form (30). It is checked that $L$ is generated by the restricted grammar of index $k$ $G=\langle V,\Sigma,M,S\rangle$, where $V=\{S,X_{1},\ldots,X_{k}\}$ and $M$ is given by the matrices
  • $(S\rightarrow X_{1}\cdots X_{k})$,
  • $(X_{1}\rightarrow\varepsilon,\ldots,X_{k}\rightarrow\varepsilon)$,
  • $(S\rightarrow u_{1}^{{\bf v}_{01}}X_{1}\cdots u_{k}^{{\bf v}_{0k}}X_{k})$, with ${{\bf v}_{0}=({\bf v}_{01}},\ldots,{\bf v}_{0k})$,
@@
Example 2 (continued). Let ${\cal Q}=(s,\varphi,P,\lambda)$ be a finite quantum automaton where $\varphi:\Sigma^{*}\longrightarrow O_{n}$ is the morphism (4) associated with ${\cal Q}$. Recall that $\mathbf{C}l{({\cal A})}$ and $\overline{{\cal A}}$ denote, respectively, the Euclidean and the Zariski closure of a set $\cal A$. The set ${\mathbf{C}l}(\varphi(L))$ is then equal to
@@
We observe that, not only, Proposition 10 provides a unitary scheme for the decidability results of [10] (i.e., Corollary 2 and Corollary 4) but it gives a non trivial extension of them. Indeed the following example provides languages, neither bounded nor context-free, defined by restricted grammars, for which the Intersection Problem is now decidable.
@@ 6.2 The case of monoidal languages
We start with the following definition that is instrumental for this section. We first recall that a context-free linear grammar is said to be minimal linear if it has a unique variable.
@@
of $k$ families of minimal linear grammars where every terminal production $\ (X\rightarrow u),\ u\in\Sigma^{*},$ is such that $u=\varepsilon$.
@@
Let $(L,Q)$ be a pair where $Q$ is a rational quantum automaton and $L$ is a language generated by a monoidal grammar $G={\cal G}_{1}\circ{\cal G}_{2}\circ\cdots\circ{\cal G}_{k}$. If the Zariski closures of the monoids of cycles of the grammars of ${\cal G}_{k}$ – i.e., the lowest level of $G$ – are irreducible, then ${\mathbf{C}l}(\varphi(L))$ is effective semialgebraic.
Remark 3
Examples of monoids generated by a finite number of matrices whose Zariski closure is irreducible can be borrowed from [16]. In Theorem 1.1 of loc.cit., the authors determine the Zariski closure of a cyclic matrix semigroup over $\mathbb{C}$. By the proof [22, Corollary 17], the above result implies that if we consider the cyclic matrix semigroup generated by a real matrix with eigenvalues greater than $1$, its Zariski closure is irreducible over $\mathbb{R}$. Similarly, Example 3.8 from [16] can be adapted to produce an example of a matrix semigroup $M$, generated by two matrices, such that the Zariski closure of $M$ is irreducible over $\mathbb{R}$.
We will do the proof of Proposition 11 for $k\leq 2$, i.e., for a monoidal grammar $G={\cal G}_{1}\circ{\cal G}_{2}$ of index $2$, since, by (19), the general case is treated with the same argument. The proof is technically structured in the following lemmata. Also we find useful to recall a notation introduced in Section 2.4.
@@
Let $G={\cal G}_{1}$ be a monoidal grammar of index $1$ and let $L=L(G)$. Under the assumption of Proposition 11, there exist a set $H\in({\cal H})$ and a regular map $\Psi$ such that ${\mathbf{C}l}(\varphi(L))=\Psi(H).$ Moreover ${\mathbf{C}l}(\varphi(L))$ is effective semialgebraic and contains $I$.
@@
Set $G=\langle V,\Sigma,P,S\rangle$ and observe that $V=\{S\}$, so that $G$ has a sole monoid of cycles $M_{S}$. Taking into account (25) and the form of the terminal productions of $G$, one has ${\mathbf{C}l}(\varphi(L_{S}))=\Psi({\mathbf{C}l}(M_{S}))$, where $\Psi(X\oplus Y):=XY^{\scriptstyle T}.$ From this and the fact that, by hypothesis, ${\mathbf{C}l}(M_{S})$ is irreducible, the claim follows. ∎∎
Now assume $G={\cal G}_{1}\circ{\cal G}_{2}$ and let ${\cal G}_{1}=\langle V_{1},\Sigma_{1},P_{1},S\rangle$ be the first grammar of the composition $G$. Set $V_{1}=\{S\}$. In the sequel, for every $A\in\Sigma_{1}$, we denote by $L_{A}$ the language generated by the grammar $G_{A}\in{\cal G}_{2}$. Let $\Sigma_{1}^{*}\times\Sigma_{1}^{*}$ be the multiplicative monoid of binary relations on $\Sigma_{1}^{*}$ and let $\Pi(O_{n}\oplus O_{n})$ be the multiplicative monoid of subsets of $O_{n}\oplus O_{n}$. Define now the morphism between the two monoids
@@
There exists an effective computable finite subset $\cal R$ of $\ \Sigma_{1}^{*}\times\Sigma_{1}^{*}$ such that $M_{S}=\zeta({\cal R}^{*})$.
@@
Now observe that, by (36), $(\alpha,\beta)\in C_{S}$ so that $(\alpha,\beta^{\sim})\in\widehat{C}_{S}$. By (38) the latter implies $\varphi(u)\oplus\varphi(v)^{\scriptstyle T}\in\zeta(\widehat{C}_{S})$. Hence $M_{S}\subseteq\zeta(\widehat{C}_{S})$. The other inclusion is checked similarly. Finally observe that a finite set $\cal R$ of binary relations on $\Sigma_{1}^{*}$ generating $\widehat{C}_{S}$ as a monoid, can be effectively computed from the productions of ${\cal G}_{1}$. The lemma is proved. ∎∎
@@
${\mathbf{C}l}(M_{S})\in({\cal H}).$ In particular, ${\mathbf{C}l}(M_{S})$ is effective algebraic.